# How Sparse Can We Prune A Deep Network: A Fundamental Limit Perspective

### v3

@@ Abstract
Network pruning is a commonly used measure to alleviate the storage and computational burden of deep neural networks. However, the fundamental limit of network pruning is still lacking. To close the gap, in this work we’ll take a first-principles approach, i.e. we’ll directly impose the sparsity constraint on the loss function and leverage the framework of statistical dimension in convex geometry, thus we’re able to characterize the sharp phase transition point, i.e. the fundamental limit of the pruning ratio. Through this limit, we’re able to identify two key factors that determine the pruning ratio limit, namely, weight magnitude and network sharpness. Generally speaking, the flatter the loss landscape or the smaller the weight magnitude, the smaller pruning ratio. Moreover, we provide efficient countermeasures to address the challenges in the computation of the pruning limit, which involves accurate spectrum estimation of a large-scale and non-positive Hessian matrix. Moreover, through the lens of the pruning ratio threshold, we can provide rigorous interpretations on several heuristics in existing pruning algorithms. Extensive experiments are performed that demonstrate that our theoretical pruning ratio threshold coincides very well with the experiments. All codes are available at: https://github.com/QiaozheZhang/Global-One-shot-Pruning
@@ 1.1 Related Work
Pruning Methods: Unstructured pruning involves removing unimportant weights without adhering to some structural constraints. Typical methods in this class include: [13] presented the train-prune-retrain method, which reduces the storage and computation of neural networks by learning only the significant connections. [37; 38] employed the energy consumption of each layer to determine the pruning order and developed latency tables to identify the layers that should be pruned. [11] proposed dynamic network surgery, which reduced network complexity significantly by pruning connections in real time. [6] proposed pruning by iteratively removing part of the small weights, and based on Frankle’s iterative pruning[6], [28] introduced $l_{2}$-norm to constrain the magnitude of unimportant parameters during iterative training. To the best of our knowledge, there are still no systematic study on the fundamental limit of pruning from the theoretical perspective.
Understanding Neural Networks through Convex Geometry: Convex Geometry is a powerful tool for studying high-dimensional statistical inference [5] and learning, in specific, the neural networks, considering their high-dimensional nature. In this regard, [5] addresses the recovery threshold of the linear inverse problem – a classical statistical inference problem. In the contrast, the limit of pruning in deep neural networks is a statistical learning problem, the methods and results in [5] just cannot be simply translated to the pruning problem. [19] studied the training dimension threshold of the network from a geometric point of view by utilizing the Gordon’s Escape theorem[10], which shows that the network can be trained successfully with less degrees of freedom (DoF) in affine subspace, but the burn-in affine subspace needs a good starting point and also the lottery subspace is greatly affected by the principal components of the entire training trajectory. Therefore, essentially the DoF result in [19] provides limited knowledge about the pruning ratio threshold, which is exactly the main subject of our work. [25] studied the Lottery Tickets Hypothesis (LTH)[6] by directly applying the above DoF results in [19] to demonstrate that iteration is needed in LTH and that pruning is impacted by the eigenvalues of the loss landscape, ignoring the impact of weights magnitude. The lower bound of pruning ratio is briefly mentioned in their work, however, their predicted lower bound does not match the actual value well, in some cases even with big gap (The main reason is the spectrum estimation error in their adopted algorithm). Besides, the above works are based on Gordon’s Escape theorem [10], thus providing only the lower bound (necessary condition), and there is no discussion on the upper bound. Moreover, rigorous analysis as well as the computational issues regarding the pruning limit are all lacking. Furthermore, regarding the network sharpness, we demonstrate rigorously how it is related to the pruning lower bound, which is not touched in [25].
@@ 2 Problem Setup & Key Notions
To explore the fundamental limit of network pruning, we’ll take the first principles approach. In specific, we directly impose the sparsity constraint on the original loss function, thus the feasibility of pruning can be reduced to determining whether two sets, i.e. the sublevel set (determined by the Hessian matrix of the loss function) and the $k$-sparse set intersects. Through this framework, we’re able to leverage powerful tools in high-dimensional convex geometry, such as statistical dimension [1], Gaussian width [34] and Approximate Kinematics Formula [1].
@@ 3.1.1 Network Pruning As Set Intersection
It’s worth mentioning that this lower bound has been provided in [25] by utilizing the Gordon’s Escape Theorem[10]. The main difference between our work and [10] lies in that Gordon’s Escape Theorem is not strong enough to provide the upper bound (sufficient condition) of pruning ratio, while the Approximate Kinematic Formula we employ does.
@@ 3.1.2 Gaussain Width of the Ellipsoid
For a ellipsoid $S(\epsilon)$ defined by : $S(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf H}{\bf w}\leq\epsilon\}$, where ${\bf H}\in\mathbb{R}^{D\times D}$ is a positive definite matrix, its Gaussian width is given by:
@@
The proof of Lemma 3.3 is in Appendix D.1. The Gaussian width of an ellipsoid has been provided in [19] as in the interval $[(\sqrt{\frac{2}{\pi}}(\sum_{i=1}^{D}r_{i}^{2})^{1/2},(\sum_{i=1}^{D}r_{i}^{2})^{1/2}]$, in contrast we sharpen the estimation of Gaussian width to a point $(\sum_{i=1}^{D}r_{i}^{2})^{1/2}$. For the settings which involve projection, the squared radius $r_{i}^{2}$ should be modified to $\frac{r_{i}^{2}}{\|{\bf w}^{*}-{\bf w}^{k}\|_{2}^{2}+r_{i}^{2}}$ [19]. Therefore, the Gaussian width of projected $S(\epsilon)$ defined in Eq.(12) equals:
@@ 3.1.4 Pruning Ratio vs Magnitude & Sharpness
It is evident from Eq.(15) that for a given trained network (whose spectrum of Hessian matrix is fixed), to minimize the lower bound of pruning ratio, we just need to minimize $\|{\bf w}^{*}-{\bf w}^{k}\|_{2}$, i.e. the sum of magnitudes of the pruned parameters. Therefore, the commonly-used magnitude-based pruning algorithms get justified. Moreover, it also inspires us to employ the one-shot magnitude pruning algorithm as detailed in Section 4, whose performance proves to be better than other existing algorithms, to the best of our knowledge.
Besides the above-discussed magnitude of the pruned sub-vector, we also identify another important factor that determines the pruning ratio, i.e. the network sharpness, which describes the sharpness of the loss landscape around the minima, as defined by the trace of Hessian matrix, namely $\mathrm{Tr}({\bf H})$ ([26] and [9], network flatness is the opposite of network sharpness; as sharpness increases, flatness decreases, and vice versa.).
@@ 3.2 Upper Bound of Pruning Ratio
In order to prove the existence of a $k$-sparse weight vector ${\bf w}^{k}$, we just need to show that the all-zero vector is in $S({\bf w}^{{}^{\prime}},\epsilon)$. To that end, we’ll take advantage of the sufficient condition of the approximate kinematics formula (Theorem 2.4) to show that it suffices to render the statistical dimension of the projected cone of $S({\bf w}^{{}^{\prime}},\epsilon)$ being full, i.e. ${D-k}$. Thus we can obtain the upper bound of the number of unpruned parameters, i.e. $k$.
@@ 3.3 Fundamental Limit of Pruning Ratios
It is easy to notice that the upper bound and lower bound are nearly identical in the form, in fact, as we’ll elaborate in the following, they are also quite close in the value. In other words, we are able to obtain a sharp characterization of the fundamental limit of pruning ratio. Meanwhile, it is worthwhile noting that the pruning limit depends on the magnitude of the final weights, which is largely determined by the weight initialization. Therefore, we still need to explore whether the magnitude of final weights are dependent on the initialization values. In the appendix, we report that once the data, network architecture, and training method are fixed, the distribution of trained network weights remains nearly insensitive to the initializations. Thus we get an affirmative answer about the above question.
@@ 4 Achievable Scheme & Computational Issues
Thus far we have established the lower bound and upper bound of the pruning ratio by leveraging the Approximate Kinematic Formula in convex geometry [1]. To proceed, we will demonstrate that our obtained bounds are tight in the sense that we can devise an achievable pruning algorithm whose corresponding upper bound is quite close to the lowest possible value of the lower bound. As argued in Corollary 3.4, the magnitude pruning, which removes all the smallest $D-k$ weights, will result in the lowest pruning ratio lower bound. Inspired by this result, we’ll focus on the magnitude pruning methods in order to approach the lower bound in the sequel.
@@ 4.1 Computational Challenges & Countermeasures
[…]
@@ 4.2 Achievable Scheme: $l_{1}$ Regularization & One-shot Magnitude Pruning
After training with this relaxed objective, the network weights will be pruned based on magnitudes one time, rather than in an iterative way as in [13; 6; 28]. The performance of the above described pruning scheme (termed as "$l_{1}$ regularized one-shot magnitude pruning" and abbreviated as "LOMP") can be found in Table 10 in Appendix. The above stated "zero-dominating" property due to $l_{1}$ regularization gets supported in Figure 2(b), where it can be seen that the majority of weights are indeed extremely small.
@@
The above "zero-domination" property turns out to be of critical value for our proposed pruning scheme to nearly achieve the limit (lower bound) of the pruning ratio. Fig. 2(c) illustrates the curve $\|{\bf w}^{2}\|_{2}^{2}$, i.e. the $l_{2}$ norm of the $D-k$ smallest weights, w.r.t. $k/D$. The vertical line therein represents $\rho_{L}$, the lower bound of the pruning ratio predicted in Section 3.1. When $k=D\rho_{L}$, the curve and the line will intersect as shown in Figure 2(c). Mathematically, the upper bound for the pruning ratio can be approximated as follows:
@@ 5.2 Prediction Performance
CIFAR10 FC5 2.1$\pm$0.25 1.7$\pm$0.12 -0.40$\pm$0.35
FC12 1.0$\pm$0.30 0.8$\pm$0.06 -0.24$\pm$0.33
AlexNet 0.9$\pm$0.00 0.8$\pm$0.08 -0.14$\pm$0.08
VGG16 0.8$\pm$0.06 0.8$\pm$0.08 0.04$\pm$0.08
CIFAR100 ResNet18 1.5$\pm$0.05 2.0$\pm$0.13 0.54$\pm$0.15
ResNet50 1.9$\pm$0.05 2.1$\pm$0.16 0.28$\pm$0.19
TinyImagenet ResNet18 3.9$\pm$0.82 4.3$\pm$0.38 0.46$\pm$0.71
ResNet50 2.6$\pm$0.24 2.9$\pm$0.33 0.36$\pm$0.10
@@ 7 Conclusion
In this paper we explore the fundamental limit of pruning ratio of deep networks by taking the first principles approach and leveraging the framework of convex geometry, thus we can reduce the pruning limit problem to the sets intersection problem and by taking advantage of the powerful Approximate Kinematic Formula, we can sharply characterize the fundamental limit of network pruning ratio. This fundamental limit convey a key message as follows: the network pruning limit is mainly determined by the weight magnitude and the network sharpness. These two guidelines can provide intuitive explanations of several heuristics in existing pruning algorithms. Moreover, to address the challenges in computing the involved Gaussian width, we develop an improved spectrum estimation for large-scale and non-positive Hessian matrices. Experiments demonstrate the almost perfect agreement between our theoretical results and the experimental ones.
@@
We demonstrate that when the magnitude of the removed weights is small, the upper bound of the pruning ratio nearly coincides with the lower bound. However, not all training strategies result in small removed weights, and in such cases, the upper and lower bounds do not coincide. Due to the intrinsic challenge of empirical risk minimization (ERM), specifically, overfitting, some form of regularization is necessary. As a result, it becomes difficult to examine the pruning limits of the network in its purest form(without regularization). To make progress, we turn to the most direct form of regularization, aiming to find the sparsest solution, namely $l_{0}$ regularization and its convex relaxation, $l_{1}$ regularization. The remaining question is under what conditions $l_{1}$ regularization will yield the same pruning limits as $l_{0}$ regularization and is still under investigation.
@@ B.2 Model
conv4 $128\times 128\times 3\times 3$ 1 1
@@
conv5 $128\times 256\times 3\times 3$ 1 1
@@ C.2 Gaussian Width of the Deformed Ellipsoid
Consider a well-trained neural network with weights ${\bf w}$, whose loss function defined by ${\bf w}$ has a Hessian matrix ${\bf H}$. Due to the non-positive definiteness of ${\bf H}$, there exist negative eigenvalues. Let the eigenvalue decomposition of ${\bf H}$ be ${\bf H}={\bf v}^{T}{\bf\Sigma v}$, where ${\bf\Sigma}$ is a diagonal matrix of eigenvalues. Let ${\bf D}={\bf v}^{T}{\bf|\Sigma|v}$, where ||$\cdot$|| ⋅ | means absolute operation. the geometric objects defined by H and D are $S(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf H}{\bf w}\leq\epsilon\}$ and $\hat{S}(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf D}{\bf w}\leq\epsilon\}$, then:
@@ D.1 Gaussian Width of the Quadratic Well
$={\bf x^{T}U\Sigma U^{T}x}$ Eigenvalue Decomposition of $\quad\quad\quad\quad\text{Eigenvalue Decomposition of }{\bf Q}:\ {\bf Q=U\Sigma U^{T}}.$
[…]
@@
$\mathbb{E}\sqrt{f({\bf x})}$ $\approx\mathbb{E}\sqrt{\mathbb{E}f({\bf x})+4\sqrt{\mathrm{ln}\epsilon_{1}\mathrm{ln}\epsilon_{2}}\lambda_{max}}$ Concentration of $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{Concentration of }f({\bf x}).$
[…]
@@
A (closed) $ball\ B(c,r)$ (in $\mathbb{R}^{D}$) centered at $c\in\mathbb{R}^{D}$ with radius $r$ is the set
@@
The set $B(0,1)$ is called the $unit\ ball$. An ellipsoid is just an affine transformation of a ball.
@@
. An $ellipsoid\ S$ centered at the origin is the image $L(B(0,1))$ of the unit ball under an $invertible$ linear transformation $L:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. An ellipsoid centered at a general point $c\in\mathbb{R}^{D}$ is just the translate $c+S$ of some ellipsoid $S$ centered at the origin.
@@
$=\mathbb{E}\lVert{\bf L^{T}g}\rVert_{2}\quad\quad\quad\quad\quad\quad\quad\quad\quad$ Definition of Ball.
$=\mathbb{E}\sqrt{{\bf g^{T}LL^{T}g}}\quad\quad\quad\quad\quad$ $\lVert{\bf g}\rVert_{2}=\sqrt{\bf g^{T}g},\ \text{where}\ {\bf g}\in\mathbb{R}^{D\times 1}.$
@@
$=\sqrt{2\epsilon\mathrm{Tr}({\bf H}^{-1})}\quad\quad\quad\quad\quad\quad\quad$ Invariance of Gaussian under rotation.

### v5

@@ Abstract
Network pruning is a commonly used measure to alleviate the storage and computational burden of deep neural networks. However, the fundamental limit of network pruning is still lacking. To close the gap, in this work we’ll take a first-principles approach, i.e. we’ll directly impose the sparsity constraint on the loss function and leverage the framework of statistical dimension in convex geometry, thus enabling us to characterize the sharp phase transition point, which can be regarded as the fundamental limit of the pruning ratio. Through this limit, we’re able to identify two key factors that determine the pruning ratio limit, namely, weight magnitude and network sharpness. Generally speaking, the flatter the loss landscape or the smaller the weight magnitude, the smaller pruning ratio. Moreover, we provide efficient countermeasures to address the challenges in the computation of the pruning limit, which mainly involves the accurate spectrum estimation of a large-scale and non-positive Hessian matrix. Moreover, through the lens of the pruning ratio threshold, we can also provide rigorous interpretations on several heuristics in existing pruning algorithms. Extensive experiments are performed which demonstrate that our theoretical pruning ratio threshold coincides very well with the experiments. All codes are available at: https://github.com/QiaozheZhang/Global-One-shot-Pruning
@@ 1.1 Related Work
Pruning Methods: Unstructured pruning involves removing unimportant weights without adhering to some structural constraints. Typical methods in this class include: [13] presented the train-prune-retrain method, which reduces the storage and computation of neural networks by learning only the significant connections. [37; 38] employed the energy consumption of each layer as the metric to determine the pruning order and developed latency tables to identify the layers that should be pruned. [11] proposed dynamic network surgery, which reduced network complexity significantly by pruning connections in real time. [6] proposed pruning by iteratively removing part of the small weights, and based on Frankle’s iterative pruning[6], [28] introduced $l_{2}$-norm to constrain the magnitude of unimportant parameters during iterative training. To the best of our knowledge, there is still no systematic study on the fundamental limit of pruning from the theoretical perspective.
Understanding Neural Networks through Convex Geometry: Convex Geometry is a powerful tool for characterizing the performance limit of high-dimensional statistical inference [5] and learning problems. For statistical inference, [5] pioneered to employ the convex geometry to study the recovery threshold of the classical linear inverse problem. For statistical learning, [19] studied the training dimension threshold of the network from a geometric point of view by utilizing the Gordon’s Escape theorem[10], which shows that the network can be trained with less degrees of freedom (DoF) than the network size in the affine subspace. The most relevant work to ours is [25], which studied the Lottery Tickets Hypothesis (LTH)[6] by applying the above DoF results in [19] to demonstrate that iteration is needed in LTH and that pruning is impacted by the eigenvalues of the loss landscape. The main difference between [25] and ours are as follows: 1) The lower bound of the pruning ratio is only briefly mentioned in [25], and their predicted lower bound does not match the actual value well, in some cases even with quite big gap (the main reason lies in the spectrum estimation error in their algorithm). 2) The core results in [25] are based on Gordon’s Escape theorem [10], which can only provide the lower bound (necessary condition). 3) Rigorous analysis as well as the computational issues regarding the pruning limit are lacking in [25]. In contrast, all the above issues are addressed in this paper.
@@ 2 Problem Setup & Key Notions
To explore the fundamental limit of network pruning, we’ll take the first principles approach. In specific, we directly impose the sparsity constraint on the original loss function, thus the feasibility of pruning can be reduced to determining whether two sets, i.e. the sublevel set (determined by the Hessian matrix of the loss function) and the $k$-sparse set intersects. Through this framework, we’re able to leverage tools in high-dimensional convex geometry, such as statistical dimension [1], Gaussian width [34] and Approximate Kinematics Formula [1].
@@ 3.1.1 Network Pruning As Set Intersection
It’s worth mentioning that this lower bound has been provided in [25] by utilizing the Gordon’s Escape Theorem[10]. The main difference between our work and [10] lies in that Gordon’s Escape Theorem is not strong enough to provide the upper bound (sufficient condition) of the pruning ratio, while the Approximate Kinematic Formula we employ does.
@@ 3.1.2 Gaussain Width of the Ellipsoid
For an ellipsoid $S(\epsilon)$ defined by : $S(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf H}{\bf w}\leq\epsilon\}$, where ${\bf H}\in\mathbb{R}^{D\times D}$ is a positive definite matrix, its Gaussian width is given by:
@@
The proof of Lemma 3.3 is in Appendix D.1. The Gaussian width of an ellipsoid has been provided in [19] as in the interval $(\sqrt{\frac{2}{\pi}}(\sum_{i=1}^{D}r_{i}^{2})^{1/2},(\sum_{i=1}^{D}r_{i}^{2})^{1/2}]$, in contrast we sharpen the estimation of Gaussian width to a point $(\sum_{i=1}^{D}r_{i}^{2})^{1/2}$. For the settings which involve projection, the squared radius $r_{i}^{2}$ should be modified to $\frac{r_{i}^{2}}{\|{\bf w}^{*}-{\bf w}^{k}\|_{2}^{2}+r_{i}^{2}}$ [19]. Therefore, the Gaussian width of projected $S(\epsilon)$ defined in Eq.(12) equals:
@@ 3.1.4 Pruning Ratio vs Magnitude & Sharpness
It is evident from Eq.(15) that for a given trained network (whose spectrum of the Hessian matrix is fixed), to minimize the lower bound of the pruning ratio, we just need to minimize $\|{\bf w}^{*}-{\bf w}^{k}\|_{2}$, i.e. the sum of magnitudes of the pruned parameters. Therefore, the commonly-used magnitude-based pruning algorithms get justified. Moreover, it also inspires us to employ the one-shot magnitude pruning algorithm as detailed in Section 4, whose performance proves to be better than other existing algorithms, to the best of our knowledge.
Besides the above-discussed magnitude of the pruned sub-vector, we also identify another important factor that determines the pruning ratio, i.e. the network sharpness, which describes the sharpness of the loss landscape around the minima, as defined by the trace of the Hessian matrix, namely $\mathrm{Tr}({\bf H})$ ([26] and [9], network flatness is the opposite of network sharpness; as sharpness increases, flatness decreases, and vice versa.).
@@ 3.2 Upper Bound of Pruning Ratio
In order to prove the existence of a $k$-sparse weight vector ${\bf w}^{k}$, we just need to show that the all-zero vector is in $S({\bf w}^{{}^{\prime}},\epsilon)$. To that end, we’ll take advantage of the sufficient condition of the approximate kinematics formula (Theorem 2.4) to show that it suffices to render the statistical dimension of the projected cone of $S({\bf w}^{{}^{\prime}},\epsilon)$ being full, i.e. Dk{D-k}. Thus we can obtain the upper bound of the number of unpruned parameters, i.e. $k$.
@@ 3.3 Fundamental Limit of Pruning Ratios
It is easy to notice that the upper bound and lower bound are of nearly identical form. In fact, as we’ll elaborate in the following, they are also of quite close value, which implies that we are able to obtain a sharp characterization of the fundamental limit of pruning ratio. Meanwhile, it is worthwhile noting that the pruning limit depends on the magnitude of the final weights, which might be significantly impacted by the weight initialization. Therefore, we still need to explore whether the magnitude of final weights is dependent on the initialization values. In the appendix, we’ll demonstrate that once the data, network architecture and training method are fixed, the distribution of trained network weights remains nearly insensitive to the initializations, thus yielding an affirmative answer about the above question.
@@ 4 Achievable Scheme & Computational Issues
Thus far we have established the lower bound and upper bound of the pruning ratio by leveraging the Approximate Kinematic Formula in convex geometry [1]. To proceed, we will demonstrate that our obtained bounds are tight in the sense that we can devise an achievable pruning algorithm whose corresponding upper bound is quite close to the lowest possible value of the lower bound. As argued in Corollary 3.4, the magnitude pruning, which removes all the smallest DkD-k weights, will result in the lowest pruning ratio lower bound. Inspired by this result, we’ll focus on the magnitude pruning methods in order to approach the lower bound in the sequel.
@@ 4.1 Computational Challenges & Countermeasures
Input: Hermitian matrix $\bf A$} of size $n\times n$, Lanczos iterations $m$, ESD computation iterations $l$, gaussian kernel $f$ and variance $\sigma^{2}$, sampling numss and row $l_{1}$ norm threshold $\epsilon$.
@@ 4.2 Achievable Scheme: $l_{1}$ Regularization & One-shot Magnitude Pruning
After training with this relaxed objective, the network weights will be pruned based on magnitudes one time, rather than in an iterative way as in [13; 6; 28]. The performance of the above described pruning scheme (termed as $l_{1}$ regularized one-shot magnitude pruning" and abbreviated as "LOMP") can be found in Table 10 in Appendix. The above stated "zero-dominating" property due to $l_{1}$ regularization gets supported in Figure 2(b), where it can be seen that the majority of weights are indeed extremely small.
@@
The above "zero-domination" property turns out to be of critical value for our proposed pruning scheme to nearly achieve the limit (lower bound) of the pruning ratio. Fig. 2(c) illustrates the curve $\|{\bf w}^{2}\|_{2}^{2}$, i.e. the $l_{2}$ norm of the DkD-k smallest weights, w.r.t. $k/D$. The vertical line therein represents $\rho_{L}$, the lower bound of the pruning ratio predicted in Section 3.1. When $k=D\rho_{L}$, the curve and the line will intersect as shown in Figure 2(c). Mathematically, the upper bound for the pruning ratio can be approximated as follows:
@@ 5.2 Prediction Performance
[…]
@@ 7 Conclusion
In this paper we explore the fundamental limit of pruning ratio of deep networks by taking the first principles approach and leveraging the framework of convex geometry. Specifically, we reduce the pruning limit problem to the sets intersection problem, and by taking advantage of the powerful Approximate Kinematic Formula, we are able to sharply characterize the fundamental limit of the network pruning ratio. This fundamental limit conveys a key message as follows: the network pruning limit is mainly determined by the weight magnitude and the network sharpness. These two guidelines can provide intuitive explanations of several heuristics in existing pruning algorithms. Moreover, to address the challenges in computing the involved Gaussian width, we develop an improved spectrum estimation for large-scale and non-positive Hessian matrices. Experiments demonstrate the almost perfect agreement between our theoretical results and the experimental ones.
@@
In this paper, the (almost) coincidence of the upper bound and lower bound of the pruning ratio depends on the condition that the removed weights are of quite small value, which is enabled by the $l_{1}$ regularization we employed. Therefore, it is important to demonstrate whether the $l_{1}$ regularized training is optimal or nearly optimal in the sense of obtaining the smallest sub-network without performance degradation for the original learning problem.
@@ B.2 Model
conv4 $28\times 128\times 3\times 3 1 1$
@@
conv5 $28\times 256\times 3\times 3 1 1$
@@ C.2 Gaussian Width of the Deformed Ellipsoid
Consider a well-trained neural network with weights $\bf w$}, whose loss function defined by $\bf w$} has a Hessian matrix $\bf H$}. Due to the non-positive definiteness of $\bf H$}, there exist negative eigenvalues. Let the eigenvalue decomposition of $\bf H$} be ${\bf H}={\bf v}^{T}{\bf\Sigma v}$, where $\bf\Sigma$} is a diagonal matrix of eigenvalues. Let ${\bf D}={\bf v}^{T}{\bf|\Sigma|v}$, where $|\cdot|$ means absolute operation. the geometric objects defined by H and D are $S(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf H}{\bf w}\leq\epsilon\}$ and $\hat{S}(\epsilon):=\{{\bf w}\in\mathbb{R}^{D}:\frac{1}{2}{\bf w}^{T}{\bf D}{\bf w}\leq\epsilon\}$, then:
@@ D.1 Gaussian Width of the Quadratic Well
$={\bf x^{T}U\Sigma U^{T}x}$ $\quad\quad\quad\quad\text{Eigenvalue Decomposition of }{\bf Q}:\ {\bf Q=U\Sigma U^{T}}.$
$=\sum_{i=1}^{n}\lambda_{i}x_{i}^{2}\quad\text{w. p. almost 1.}$ Invariance of Gaussian under rotation.
@@
$\mathbb{E}\sqrt{f({\bf x})}$ $\approx\mathbb{E}\sqrt{\mathbb{E}f({\bf x})+4\sqrt{\mathrm{ln}\epsilon_{1}\mathrm{ln}\epsilon_{2}}\lambda_{max}}$ $\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{Concentration of }f({\bf x}).$
$\approx\sqrt{\mathbb{E}f({\bf x})}+\frac{2\sqrt{\mathrm{ln}\epsilon_{1}\mathrm{ln}\epsilon_{2}}\lambda_{max}}{\sqrt{\mathbb{E}f({\bf x})}}$ $\text{ Taylor Expansion}.$
@@
A (closed) ballB(c,r)ball\ B(c,r) (in $\mathbb{R}^{D}$) centered at $c\in\mathbb{R}^{D}$ with radius $r$ is the set
@@
The set $B(0,1)$ is called the unitballunit\ ball. An ellipsoid is just an affine transformation of a ball.
@@
. An ellipsoidSellipsoid\ S centered at the origin is the image $L(B(0,1))$ of the unit ball under an invertibleinvertible linear transformation $L:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. An ellipsoid centered at a general point $c\in\mathbb{R}^{D}$ is just the translate $c+S$ of some ellipsoid $S$ centered at the origin.
@@
$=\mathbb{E}\lVert{\bf L^{T}g}\rVert_{2}\quad\quad\quad\quad\quad\quad\quad\quad\quad\ \ \$Definition of Ball.
$=\mathbb{E}\sqrt{{\bf g^{T}LL^{T}g}}\quad\quad\quad\quad\quad\ \ \$$\lVert{\bf g}\rVert_{2}=\sqrt{\bf g^{T}g},\ \text{where}\ {\bf g}\in\mathbb{R}^{D\times 1}.$
@@
$=\sqrt{2\epsilon\mathrm{Tr}({\bf H}^{-1})}\quad\quad\quad\quad\quad\quad\quad\ \$Invariance of Gaussian under rotation.